OpenStudy (anonymous):

A hot-air balloon rises 90 feet in its first minute of flight. In each succeeding minute, it rises only 90% as far as it did during the preceeding minute. What is the final height of the balloon when it stops rising?

OpenStudy (anonymous):

The height for the \(n\)-th minute is modeled by a geometric sequence with ratio \(\dfrac{9}{10}\): \[\large\text{height}=h_n=90\left(\frac{9}{10}\right)^{n-1}\] The reason the exponent is \(n-1\) is that in the first minute (after 60 seconds, presumably), the height is 90, so when \(n=1\) you have \(h_1=90\left(\dfrac{9}{10}\right)^0=90\). The cumulative height (as \(n\to\infty\)) can be determined by computing the sum, \[\large\sum_{n=1}^\infty 90\left(\frac{9}{10}\right)^{n-1}\] Fortunately for you, there's a formula for such a series: \[\large\sum_{n=1}^\infty ar^{n-1}=\frac{a}{1-r}\]